Problem 3-20 (Algo) Plantwide versus Multiple Predetermined Overhead Rates: Service Industry [LO3-1, LO3-2, LO3-3, LO3-4]

McCullough Hospital uses a job-order costing system to assign costs to its patients. Its direct materials include a variety of items such as pharmaceutical drugs, heart valves, artificial hips, and pacemakers. Its direct labor costs (e.g., surgeons, anesthesiologists, radiologists, and nurses) associated with specific surgical procedures and tests are traced to individual patients. All other costs, such as depreciation of medical equipment, insurance, utilities, incidental medical supplies, and the labor costs associated with around-the-clock monitoring of patients are treated as overhead costs.

Historically, McCullough has used one predetermined overhead rate based on the number of patient-days (each night that a patient spends in the hospital counts as one patient-day) to allocate overhead costs to patients. Recently a member of the hospital’s accounting staff has suggested using two predetermined overhead rates (allocated based on the number of patient-days) to improve the accuracy of the costs allocated to patients. The first overhead rate would include all overhead costs within the Intensive Care Unit (ICU) and the second overhead rate would include all Other overhead costs. Information pertaining to the hospital’s estimated number of patient-days, its estimated overhead costs, and two of its patients—Patient A and Patient B—is provided below:

ICU Other Total
Estimated number of patient-days 2,500 22,500 25,000
Estimated fixed overhead cost $ 4,122,500 $ 22,005,000 $ 26,127,500
Estimated variable overhead cost per patient-day $ 286 $ 117
Patient A Patient B
Direct materials $ 5,000 $ 6,700
Direct labor $ 26,250 $ 37,000
Total number of patient-days (including ICU) 14 19
Number of patient-days spent in ICU 0 7

Required:

1. Assuming McCullough uses only one predetermined overhead rate, calculate:

a. The predetermined overhead rate.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

2. Assuming McCullough calculates two overhead rates as recommended by the staff accountant, calculate:

a. The ICU and Other overhead rates.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

(Round "Predetermined overhead rate" to 2 decimal places. Round your intermediate calculations and final answers to the nearest dollar amount.)
Answer #1
1a
Estimated fixed overhead cost 26127500
Estimated variable overhead cost 3347500 =(2500*286)+(22500*117)
Estimated total overhead cost 29475000
Divide by Total Estimated number of patient-days 25000
The predetermined overhead rate 1179 per patient day
b
Patient A
Direct materials 5000
Direct labor 26250
Applied overhead 16506 =14*1179
Total cost for Patient A 47756
Patient B
Direct materials 6700
Direct labor 37000
Applied overhead 22401 =19*1179
Total cost for Patient B 66101
2a
Estimated fixed overhead cost 4122500
Estimated variable overhead cost 715000 =2500*286
Estimated total overhead cost 4837500
Divide by Estimated number of patient-days 2500
Predetermined ICU overhead rate 1935 per patient day
Estimated fixed overhead cost 22005000
Estimated variable overhead cost 2632500 =22500*117
Estimated total overhead cost 24637500
Divide by Estimated number of patient-days 22500
Predetermined Other overhead rate 1095 per patient day
b
Patient A
Direct materials 5000
Direct labor 26250
Applied overhead 15330 =14*1095
Total cost for Patient A 46580
Patient B
Direct materials 6700
Direct labor 37000
Applied overhead 26685 =(7*1935)+(12*1095)
Total cost for Patient B 70385